# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 2.3.2. Snowballing
In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within 110 primary studies to identify additional relevant studies. This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies. Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search. This phase resulted in the addition of 11 more studies, bringing the final count to 121.
2.3.3. Quality Assessment of Selected Studies
@@
We assessed the quality of the selected studies based on a set of predefined criteria, including relevance to multi-cloud containerization, clarity of research design, empirical validation (e.g., experiments, case studies), reporting quality (e.g., discussion of limitations and challenges), publication in peer-reviewed venues, and the generalizability of findings to multi-cloud environment. Each study was evaluated using a scoring system where criteria were rated as 0 (not satisfied), 0.5 (partially satisfied), or 1 (fully satisfied), with a maximum total score of 6. Studies scoring below 3 were flagged for further review and discussion but retained to ensure alignment with the study’s objectives. This quality assessment was conducted after the study selection process to ensure consistency and thoroughness across all 121 studies. Detailed results are available in the replication package (Waseem et al., 2024), specifically in the sheet titled “QA of Selected Studies”.
2.4. Data Extraction and Analysis
@@
Data Analysis: We employed descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9. For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006). Our thematic analysis process consists of the following steps:
  1. (1)
@@
Table 5. Data items to be extracted in this SMS
@@
Concerning data item D9 (Roles of Containers) and D10 (Container Implementation strategy), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
Finally, we provided a replication package (Waseem et al., 2024) with the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
3. Results
@@
3.1. Demographics, Research Aims, and Contributions
Yearly Distribution of Studies: Figure 3-a illustrates the annual distribution of studies retrieved from eight databases between January 2013 and July 2024. We found only two primary studies on containerization in multi-cloud environment that meet our criteria in 2013. The bars display yearly study counts and trends. A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest followed by more targeted research. Notably, the reduction in publication numbers in 2020 and 2021 may be partially attributed to the global disruptions caused by the COVID-19 pandemic, which affected research outputs across various fields. However, there is a notable increase in studies in 2023 and continuing into 2024. This could be due to delayed research projects getting back on track and more funding in tech sectors after the pandemic, showing a renewed focus on improving containerization in multi-cloud environment
Figure 3. Demographics of the Selected Studies, Including: a) Frequency and Categories of Publications; b) Publisher of the Research; c) Affiliations of the Authors
@@ 3.1. Demographics, Research Aims, and Contributions
Figure 4. Selected Studies: a) Thematic Classification and b) Trends Over the Years
Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions of the selected studies. The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis. The selected studies mainly focus on proposing and validating security solutions, reference architectures, cloud management frameworks, interoperability, orchestration, and tools for optimizing and deploying containerized applications in multi-cloud environment. Notably, the largest contributions come from the Security Approaches category, representing 27.25% of the total studies, which highlights the significant attention to cybersecurity & access control and cloud analysis approaches. Within this category, the Cybersecurity and Access Control subcategory alone accounts for 9.09%, emphasizing the importance of securing multi-cloud environment. The studies related to the Reference Architectures and Models for Cloud category, comprising 21.42% of the studies, focus on understanding and designing architectures specifically tailored for multi-cloud containerization. Within this category, General Cloud Architectures (4.92%) and Fault Tolerance and Self-Healing (4.13%) are key subcategories. These subcategories address critical challenges such as ensuring resilience and performance in cloud-native architectures. The studies on Optimization and Deployment Approaches and Tools category accounts for 18.16% of the studies, indicating a strong research focus on optimizing container deployment and resource scheduling. Within this category, Optimization & Scheduling Approaches (8.26%) represent the second-largest subcategory across all studies, highlighting the practical need for effective resource allocation in multi-cloud systems. Finally, Cloud Management and Interoperability Frameworks, accounting for 12.38%, primarily aim to improve resource utilization, service management, monitoring, and deployment in multi-cloud environments. This includes contributions from Resource and Service Management Frameworks, which represent 4.95% of the studies.
Table 6. Selected studies’ contributions on containerization in multi-cloud environment
@@
Reference Architectures and Models for Cloud General Cloud Architectures (6, 4.92%) (Watanabe et al., 2022; Sami and Mourad, 2020; Carpio et al., 2020; Baresi et al., 2016; Pop et al., 2016; Alonso et al., 2023a)
IoT & Service Mesh Architecture (5, 4.13%) (Gattobigio et al., 2022; Rodigari et al., 2021; Merlino et al., 2024; Hossain et al., 2023; Solayman and Qasha, 2023)
Fault Tolerance & Self-Healing (5, 4.13%) (Bisht and Kaur, 2022; Bolivar et al., 2018; Tang, 2021; Mfula and Nurminen, 2018; Nagarajan et al., 2023)
Migration & Testing Architecture (3, 2.47%) (Slominski et al., 2015; Serhiienko and Spillner, 2018; Roy and Kalafatis, 2023)
Optimization & Reliability Model (3, 2.47%) (Bhamare et al., 2015; Aral et al., 2020; Huang and Wang, 2023)
Domain-Specific Language (2, 1.65%) (Quint and Kratzke, 2018; Ferry et al., 2013)
Performance Modeling (2, 1.65%) (Barna et al., 2017; Rahman and Lama, 2019)
Security Approaches Cybersecurity & Access Control (11, 9.09%) (He et al., 2022; Sheridan et al., 2017; Rios et al., 2017; Graupner et al., 2015a; Guechi and Maamri, [n. d.])
Cloud Analysis Approaches (7, 5.78%) (Pai and Kumar, 2021; Pahl, 2015; Dimitrijevic et al., 2019; Zeck and Bouroudjian, 2017; Kiss, 2018; Gundu et al., 2020)
Security and Performance Solution (6, 4.95%) (Sabbioni et al., 2020; Wang et al., 2020; Pandey et al., 2022; Ranjbar et al., 2017)
Security Approaches (5, 4.13%) (Deng et al., 2022; Gao et al., 2017; Bélair et al., 2019; Wahab et al., 2016)
Provisioning & Restoration Approach (4, 3.30%) (Pellegrini et al., 2017; Lazuka et al., 2022; Wang et al., 2019; Kritikos et al., 2019)
Cloud Management and Interoperability Framework Resource & Service Management Framework (6, 4.95%) (Saif et al., 2023; Harwalkar et al., 2019b; Libardi et al., 2014; Pham and Pham, 2015; Sitaram et al., 2018; Barletta et al., 2024)
Monitoring & Deployment Framework (5, 4.13%) (Alhamazani et al., 2014; Trihinas et al., 2015; Fadda et al., 2019; Pham et al., 2015; Horchulhack et al., 2024; Marques et al., 2024)
Evaluation & Modeling Framework (4, 3.30%) (Ferry et al., 2014; Bou Ghantous and Gill, 2021; Tundo et al., 2024; Ramesh et al., 2023)
Interoperability Framework (2, 1.65%) (Hua et al., 2020; Challita et al., 2017)
Orchestration and Deployment Approach Benchmarking & Optimization Approach (6, 4.13%) (Almeida et al., 2014; Jha et al., 2019; Shi et al., 2020; Buyya and Barreto, 2015; Guerrero et al., 2018a; Petrovski and Gusev, 2024)
Service Management Approach (4, 3.30%) (Zhang et al., 2021; Chen et al., 2022; Gupta et al., 2021; Benmerar et al., 2023)
Simulation & Integration Approach (3, 2.47%) (Alonso et al., 2019; Mittal and Risco-Martín, 2017; del Castillo et al., 2013)
Storage & Services Solution (3, 2.47%) (Harwalkar et al., 2019a; Shen et al., 2017; Orzechowski et al., 2023)
Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach (10, 8.26%) (Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; da Silva et al., 2023; Dogani et al., 2023; Pandi et al., 2023; Sofia et al., 2023; Nguyen and Chien, 2023)
Multi-Cloud Platform (6, 4.95%) (Verdugo et al., 2017; Ouyang et al., 2022; Dreibholz et al., 2019; Janarthanan et al., 2018; Alaluna et al., 2017; Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; Santos et al., 2023)
Deployment Approach and Tool (6, 4.95%) (Cui et al., 2016; Casola et al., 2017; Tricomi et al., 2017; Di Modica et al., 2019; Liu and Qin, 2024; Tiwari and Sharma, 2023)
[…]
3.2. Container Roles and Implementation Strategies (RQ1)
We report the results related to roles and implementation strategies in the following two subsections.
3.2.1. Container Roles
@@
Table 7. Roles of containers in multi-cloud environment
@@
Container-Based Resource and Service Management Container-Based Multi-Tenancy Solution (Ramos et al., 2021)
Container-Driven Resource Sharing & Multi-Tenancy (Deng et al., 2022)
Multi-Tenant Service Management via Containers (Slominski et al., 2015)
Container-Enabled OS-Level Virtualization (Bélair et al., 2019)
Container-Based Complexity Reduction (Shi et al., 2020)
Containerized Service Hosting (Pham and Pham, 2015)
Multi-Cloud Workload Migration via Containers (Janarthanan et al., 2018; Zeck and Bouroudjian, 2017)
Seamless Workload Migration via Containers (Zeck and Bouroudjian, 2017)
Container-Based Multi-Cloud Platform Execution (Serhiienko and Spillner, 2018)
Container-Driven MicroCloud Scaling & Isolation (Baresi et al., 2016)
Container-based Application Deployment Containerized Web-App Deployment (Almeida et al., 2014)
Container-Driven App Deployment (Gattobigio et al., 2022; Pham et al., 2015)
Containerized PaaS Prototyping (Verdugo et al., 2017)
Data-Intensive App Deployment via Containers (Barna et al., 2017)
Containerized App Management (Gupta et al., 2021)
Container-Based Aneka Service Deployment (Buyya and Barreto, 2015)
Container-Driven Microservices Deployment & Scaling (Rahman and Lama, 2019)
Container-Based Multi-Cloud App Deployment & Scaling (Mfula and Nurminen, 2018)
Containerized IoT App Deployment (Bou Ghantous and Gill, 2021)
Containerization in App Development & IoT App Development & IoT Integration via Containers (Ouyang et al., 2022)
Container-Based OpenVAS & Integration of Chef (Sheridan et al., 2017)
Containers as Standardized Software Units (Al-shammari and Alwan, 2018)
Container Placement & Orchestration in Fog Computing (Sami and Mourad, 2020)
Container Deployment Speed & DevOps Integration (Lorenzo et al., 2019)
Containerizing Apps for Privacy & Compliance (Dimitrijevic et al., 2019)
Container-Driven Portability & Efficiency Improvement (Altran et al., 2022)
Preventing Vendor Lock-In via Containers (Pellegrini et al., 2017)
Container-Based Extensible App Development (Pop et al., 2016)
Performance & Efficiency via Containers Scalable App Building via Containers (Pai and Kumar, 2021)
Container-Driven SaaS Performance Isolation (Wang et al., 2020)
Cloud Performance Enhancement via Containers (Saif et al., 2023)
Evaluating Performance Impact of Containers (Rodigari et al., 2021)
Cost Efficiency of Microservices via Containers (Guerrero et al., 2018a)
Container-Based Cloud Platform Optimization (Kiss, 2018)
Container Technology Features Container Usage in Edge Computing (Wong et al., 2019)
Container-Enabled Lightweight Virtualization (Pahl, 2015)
Key Features of Container Technology (Gao et al., 2017)
Container-Based Edge Computing Virtualization (Carpio et al., 2020)
Container Leverage in Kubernetes (Lazuka et al., 2022)
Other Container Uses Container-Based Azure Data Storage (Bucur et al., 2018)
Container & VM Support in Architecture (Bolivar et al., 2018)
Blockchain-Based Multi-Cloud Reliability via Containers (Aral et al., 2020)
Evaluating Scheduling Algorithm via Containers (Tang, 2021)
Container-Based Virtual Network Composition (Alaluna et al., 2017)
[…]
3.2.2. Implementation Strategies

### v3

@@ 2.3.2 Snowballing
In Phase 2, we utilized the snowballing technique, as described in [25], to examine references within 110 primary studies to identify additional relevant studies. This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies. Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search. This phase resulted in the addition of 11 more studies, bringing the final count to 121.
2.3.3 Quality Assessment of Selected Studies
@@
We assessed the quality of the selected studies based on a set of predefined criteria, including relevance to multi-cloud containerization, clarity of research design, empirical validation (e.g., experiments, case studies), reporting quality (e.g., discussion of limitations and challenges), publication in peer-reviewed venues, and the generalizability of findings to multi-cloud environment. Each study was assessed using six predefined quality criteria: relevance to multi-cloud containerization, clarity of research design, data validity (e.g., empirical evidence), reporting quality (e.g., discussion of limitations), peer-review status, and generalizability of findings. Each criterion was scored using a 3-point scale: 1 (fully satisfied), 0.5 (partially satisfied), or 0 (not satisfied), giving a total score out of 6. The scoring was applied consistently across all studies based on observable features. For example, studies addressing both containerization and multi-cloud contexts were scored 1 for relevance; studies with explicit research methods scored higher on research design; and those lacking empirical evidence were scored lower for data validity. Scores for all individual criteria are documented in the “QA of Selected Studies” sheet in our replication package, providing transparency and allowing independent verification of the quality assessment process. Studies scoring below 3 were flagged for further review and discussion but retained to ensure alignment with the study’s objectives. This quality assessment was conducted after the study selection process to ensure consistency and thoroughness across all 121 studies. Detailed results are available in the replication package [11], specifically in the sheet titled “QA of Selected Studies”.
2.4 Data Extraction and Analysis
@@
Data Analysis: We employed descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9. For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in [26]. Our thematic analysis process consists of the following steps:
  1. 1.
@@
Table 5: Data items to be extracted in this SMS
@@
Concerning data item D9 (Roles of Containers) and D10 (Container Implementation strategy), we used the open coding and constant comparison techniques from Grounded Theory [27] to analyze the qualitative data extracted from the selected studies.
Finally, we provided a replication package [11] with the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
3 Results
@@
3.1 Demographics, Research Aims, and Contributions
Yearly Distribution of Studies: Figure 3-a illustrates the annual distribution of studies retrieved from eight databases between January 2013 and July 2024. We found only two primary studies on containerization in multi-cloud environment that meet our criteria in 2013. The bars display yearly study counts and trends. A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest followed by more targeted research. Notably, the reduction in publication numbers in 2020 and 2021 may be partially attributed to the global disruptions caused by the COVID-19 pandemic, which affected research outputs across various fields. However, there is a notable increase in studies in 2023 and continuing into 2024. This could be due to delayed research projects getting back on track and more funding in tech sectors after the pandemic, showing a renewed focus on improving containerization in multi-cloud environment.
Figure 3: Demographics of the selected studies, including: a) Frequency and Types of Publications; b) Research Publisher; c) Author’s Affiliation
@@ 3.1 Demographics, Research Aims, and Contributions
Takeaway 1: Research on containerization in multi-cloud environments has gained more attention in recent years, especially after 2022. Most of the studies are published by IEEE and come from academic authors. This shows that the topic is mainly explored in academic settings. However, there is little contribution from industry, and most studies are presented at conferences rather than in journals. This suggests a need for more collaboration between academia and industry and more detailed research through journal publications.
@@
Figure 4: Selected studies: a) Thematic Classification and b) Trends and Distribution of Themes Over the Years
Takeaway 2: The research themes show that when containerization is applied in multi-cloud environments, unique challenges emerge—such as the need for orchestration tools that can manage service consistency across diverse providers, and security mechanisms that adapt to varying cloud configurations. Studies have evolved from basic scalability and performance concerns to addressing how to efficiently monitor, deploy, and manage containerized applications across different cloud platforms using DevOps and federation-aware solutions. This highlights how multi-cloud settings demand more adaptable and interoperable container strategies than single-cloud systems.
Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions of the selected studies. The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in [11]), through thematic analysis. The selected studies mainly focus on proposing and validating security solutions, reference architectures, cloud management frameworks, interoperability, orchestration, and tools for optimizing and deploying containerized applications in multi-cloud environment. Notably, the largest contributions come from the Security Approaches category, representing 27.25% of the total studies, which highlights the significant attention to cybersecurity & access control and cloud analysis approaches. Within this category, the Cybersecurity and Access Control subcategory alone accounts for 9.09%, emphasizing the importance of securing multi-cloud environment. The studies related to the Reference Architectures and Models for Cloud category, comprising 21.42% of the studies, focus on understanding and designing architectures specifically tailored for multi-cloud containerization. Within this category, General Cloud Architectures (4.92%) and Fault Tolerance and Self-Healing (4.13%) are key subcategories. These subcategories address critical challenges such as ensuring resilience and performance in cloud-native architectures. The studies on Optimization and Deployment Approaches and Tools category accounts for 18.16% of the studies, indicating a strong research focus on optimizing container deployment and resource scheduling. Within this category, Optimization & Scheduling Approaches (8.26%) represent the second-largest subcategory across all studies, highlighting the practical need for effective resource allocation in multi-cloud systems. Finally, Cloud Management and Interoperability Frameworks, accounting for 12.38%, primarily aim to improve resource utilization, service management, monitoring, and deployment in multi-cloud environments. This includes contributions from Resource and Service Management Frameworks, which represent 4.95% of the studies.
Table 6: Selected studies’ contributions on containerization in multi-cloud environment
@@
Reference Architectures and Models for Cloud General Cloud Architectures (6, 4.92%) [28, 29, 30, 31, 32, 33]
IoT & Service Mesh Architecture (5, 4.13%) [34, 35, 36, 37, 38]
Fault Tolerance & Self-Healing (5, 4.13%) [39, 40, 41, 42, 43]
Migration & Testing Architecture (3, 2.47%) [44, 45, 46]
Optimization & Reliability Model (3, 2.47%) [47, 48, 49]
Domain-Specific Language (2, 1.65%) [50, 51]
Performance Modeling (2, 1.65%) [52, 53]
Security Approaches Cybersecurity & Access Control (11, 9.09%) [54, 55, 56, 57, 58]
Cloud Analysis Approaches (7, 5.78%) [59, 60, 61, 62, 63, 64]
Security and Performance Solution (6, 4.95%) [65, 66, 67, 68]
Security Approaches (5, 4.13%) [69, 70, 71, 72]
Provisioning & Restoration Approach (4, 3.30%) [73, 74, 75, 76]
Cloud Management and Interoperability Framework Resource & Service Management Framework (6, 4.95%) [77, 78, 79, 80, 81, 82]
Monitoring & Deployment Framework (5, 4.13%) [83, 84, 85, 86, 87, 88]
Evaluation & Modeling Framework (4, 3.30%) [89, 90, 91, 92]
Interoperability Framework (2, 1.65%) [93, 94]
Orchestration and Deployment Approach Benchmarking & Optimization Approach (6, 4.13%) [95, 96, 97, 98, 99, 100]
Service Management Approach (4, 3.30%) [101, 102, 103, 104]
Simulation & Integration Approach (3, 2.47%) [105, 106, 107]
Storage & Services Solution (3, 2.47%) [108, 109, 110]
Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach (10, 8.26%) [111, 112, 113, 114, 115, 116, 117, 118, 119]
Multi-Cloud Platform (6, 4.95%) [120, 121, 122, 123, 124, 111, 112, 113, 114, 125]
Deployment Approach and Tool (6, 4.95%) [126, 127, 128, 129, 130, 131]
3.2 Container Roles and Implementation Strategies (RQ1)
In this section, we present the roles that containers play in multi-cloud environments and the strategies used to implement them. Based on thematic analysis of the selected studies, we classified the findings into six major roles and five strategic categories. Each category is supported with examples to help readers relate to real-world use cases. The goal is to provide a holistic understanding of how containers are applied and managed across clouds. Readers can interpret the tables as a catalog of practical use scenarios and corresponding strategic techniques.
3.2.1 Container Roles
@@
Takeaway 3: The contributions of the reviewed studies highlight how multi-cloud environments introduce specific demands on containerization, particularly in areas such as cross-platform security, cloud-aware architecture design, and adaptive resource scheduling. Unlike generic solutions, these studies propose container orchestration frameworks, interoperability tools, and security models tailored to the challenges of managing distributed containerized applications across heterogeneous cloud providers.
@@
Table 7: Roles of containers in multi-cloud environment
@@
Container-Based Resource and Service Management Container-Based Multi-Tenancy Solution [132]
Container-Driven Resource Sharing & Multi-Tenancy [69]
Multi-Tenant Service Management via Containers [44]
Container-Enabled OS-Level Virtualization [71]
Container-Based Complexity Reduction [97]
Containerized Service Hosting [80]
Multi-Cloud Workload Migration via Containers [123, 62]
Seamless Workload Migration via Containers [62]
Container-Based Multi-Cloud Platform Execution [45]
Container-Driven MicroCloud Scaling & Isolation [31]
Container-based Application Deployment Containerized Web-App Deployment [95]
Container-Driven App Deployment [34, 86]
Containerized PaaS Prototyping [120]
Data-Intensive App Deployment via Containers [52]
Containerized App Management [103]
Container-Based Aneka Service Deployment [98]
Container-Driven Microservices Deployment & Scaling [53]
Container-Based Multi-Cloud App Deployment & Scaling [42]
Containerized IoT App Deployment [90]
Containerization in App Development & IoT App Development & IoT Integration via Containers [121]
Container-Based OpenVAS & Integration of Chef [55]
Containers as Standardized Software Units [112]
Container Placement & Orchestration in Fog Computing [29]
Container Deployment Speed & DevOps Integration [113]
Containerizing Apps for Privacy & Compliance [61]
Container-Driven Portability & Efficiency Improvement [114]
Preventing Vendor Lock-In via Containers [73]
Container-Based Extensible App Development [32]
Performance & Efficiency via Containers Scalable App Building via Containers [59]
Container-Driven SaaS Performance Isolation [66]
Cloud Performance Enhancement via Containers [77]
Evaluating Performance Impact of Containers [35]
Cost Efficiency of Microservices via Containers [99]
Container-Based Cloud Platform Optimization [63]
Container Technology Features Container Usage in Edge Computing [111]
Container-Enabled Lightweight Virtualization [60]
Key Features of Container Technology [70]
Container-Based Edge Computing Virtualization [30]
Container Leverage in Kubernetes [74]
Other Container Uses Container-Based Azure Data Storage [133]
Container & VM Support in Architecture [40]
Blockchain-Based Multi-Cloud Reliability via Containers [48]
Evaluating Scheduling Algorithm via Containers [41]
Container-Based Virtual Network Composition [124]
Takeaway 4: The reviewed studies show that containers take on distinct roles in multi-cloud environments, particularly in addressing challenges such as provider heterogeneity, cross-cloud orchestration, and decentralized application deployment. Containers are used for multi-tenant resource management, cross-cloud service hosting, and workload migration. They also support edge and fog computing, compliance, and vendor lock-in prevention—roles less prominent in single-cloud contexts. This shows that multi-cloud scenarios introduce unique architectural and operational needs that shape how containers are implemented.
3.2.2 Implementation Strategies